OpenStudy (anonymous):

show that the set of all vectors in R3 that are orthogonal to (1,-2,2) is a subspace of R3. (use standard inner product for R3); find a basis for the subspace;

OpenStudy (amistre64):

how do we find all the vectors that are perpendicular to a given vector?

OpenStudy (amistre64):

or, how do we know when 2 vectors are perped? what property can we use?

OpenStudy (anonymous):

To find perpendicular, do the inner product of two vector and if u get 0, then they are orthogoanl

OpenStudy (amistre64):

so lets take a set of vectors (x,y,x) and the stated vector (1,-2,2) and work the dot product that defines them to be perp. wht do we get?

OpenStudy (amistre64):

*** set of vectors (x,y,z) that is ....

OpenStudy (irishboy123):

.

OpenStudy (anonymous):

I get x-2y+2z=0, if I let y=t, z=r, then I get the vector to be (2t-2r, t, r).

OpenStudy (amistre64):

x - 2y + 2z = 0 defines all the vectors in R^3 that are perp to (1,-2,2). its also known as the equation of a plane now, what are the properties we need to check to determine its a subspace?

OpenStudy (amistre64):

contains the zero vector (the origin in in the plane), well ... 0,0,0 satisfies the equation so it checks out what else we got for subspace properties?

OpenStudy (anonymous):

Closed under addition and scalar multiplication, but I don't know how to show addition & multi

OpenStudy (amistre64):

let v1 = (a1,a2,a3) and let v2=(b1,b2,b3) be any vectors that satisfy x - 2y +2z = 0 when kv1 and jv2 for some constants k and j will also satisfy the equation (be contained in the plane)

OpenStudy (amistre64):

right?

OpenStudy (amistre64):

thats for scalar multiplication and (kv1 + jv2) for addition would have to satisfy the plane equation as well

OpenStudy (amistre64):

kv1 is any vector in the plane scaled, since v1 is defined to be in the plane then kv1 outta satisfy test it out ka1 -2(ka2) +2k(a3) = 0 k (a1 -2a2 +2a3) = 0 since we know that a1,a2,a3 is in the plane by definition, then we know that a1 -2a2 +2a3=0 k (0) = 0 for all k scalar multiples of any vector in the plane

OpenStudy (amistre64):

test for addition (ka1+jb1) -2(ka2+jb2)+2(ka3+jb3)=0 ka1+jb1 -2ka2 -2jb2 +2ka3 +2jb3 = 0 k (a1 -2a2 +2a3) +j (b1 -2b2 +2b3) = 0 we now a1,a2,a3 and b1,b2,b3 satisfy the equation by definition of being in the plane already soo a1 -2a2 +2a3=0 b1 -2b2 +2b3=0 k (0) +j (0) = 0 is true for all vector additions in the plane.

OpenStudy (amistre64):

the next step is just pulling 2 linearly independant vectors to numerically define v1 and v2 and those would form the basis of the plane

OpenStudy (anonymous):

thank you so muchhhhhhh, i got it!!!! XD